Hyrkas Corporations most recent balance sheet and income statement appear below: Balance Sheet December 31, Year 2 and Year 1 (in thousands of dollars) Year 2 Year 1 Assets Current assets: Cash Accounts receivable, net Inventory Prepaid expenses $ 155200 250 170 20 640 780 230 200 20 605 790 Total current assets Plant and equipment, net Total assets $1,395 $1,420 Liabilities and Stockholders Equity Current liabilities: $ 170 200 50 40 290 190 480 Accounts payable Accrued liabilities Notes payable, short term Total current liabilities Bonds payable Total liabilities Stockholders equity: 50 40 260 160 420 Common stock, $2 par value Additional paid-in capital Retained earnings 200 330 410 940 $1,395 $1,420 200 330 445 975 Total stockholders equity Total liabilities & stockholders equityIncome Statement For the Year Ended December 31, Year 2 (in thousands of dollars) Sales (all on account) Cost of goods sold Gross margin Selling and administrative expense Net operating income Interest expense Net income before taxes Income taxes (30%) Net income $1,210 740 470 345 125 20 105 32 $ 73 Dividends on common stock during Year 2 totaled $38 thousand. The market price of common stock at the end of Year 2 was $8.70 per share.a. Gross margin percentage b. Earnings per share C. Price-earnings rati d. Dividend payout e. Dividend yield ratio f. Return on total assets 0 ratio 0 0 0 Return on equity h. Book value per share i. Working capital i, Current ratio k. Acid-test ratio Accounts receivable turnover m. Average collection period n. |Inventory turnover o. Average sale period p. | limes interest earned ratio q. Debt-to-equity ratio days days
Answer #1
a Gross Margin percentage= Gross Margin / Sales X 100
= $470/1210*100 = 38.84%
B Earnings per share = (Net Income - Dividends on Preferred Stock*) / Amount of outstanding Shares**
= ($73-0)/$200 = $.37
*As there is no preferred stock in the question, dividends on preferred stock will be zero.
** Amount of outstanding shares means value of common stock given in balance sheet which is $ 200.
There is no increase or decrease in common stock values from year 1 to year 2, so the value of common stock. $ 200 is taken for calculation.
C Price earnings ratio= Market Value per Share / Earnings per Share
D Dividend payout ratio= Dividends paid / Net income
= $38/73 = $0.5205
E Dividend yield ratio= Annual Dividend per share* / Current stock price per share X 100
= 0.38/8.70*100 = 4.37%
*Annual dividend per share = Total dividend / Number of common shares **
                                                         = $ 38,000 / 1, 00,000
                                                         = 0.38
** Number of common shares= Amount of common stock / Value per common share
                                          = $ 200,000 / $ 2 = 1,00,000
F Return on Total Assets = Earnings before interest and taxes* / Total net assets X 100
=$125/1395 = 8.961%
*Earnings before interest and taxes = Net Income + Interest expense + Taxes
                                                                          = $ 73 + $ 20 + $ 32
                                                                          = $ 125
G Return on equity = Net Income / Stockholders’ equity X 100
= $73/985*100 = 7.41%
H Book value per share = Total common stockholders’ equity / Number of common shares*
= $985/100000 = $0.0099
*Number of common shares already calculated in solution e.
I Working capital = Total current assets – Total current liabilities
= $605000-$260000 = $345000
Note: Here the amounts are taken in thousands as answer is to be calculated in thousand.
J Current Ratio = Total current assets / Total current liabilities
= $605/260 = 2.326923 times
K Acid test (quick) ratio = (Total current assets- Prepaid expenses- inventory ) / Total current liabilities
= ($605-20-$230)/$260 = 1.37 times
L Accounts receivable turnover= Net credit sales / Average accounts receivable*
= $1210/240 = 5.04 times
*Average accounts receivable= (Account receivable Year 1 + Account receivable Year 2) / 2
= ($250+230)/2 = $240
M Average collection period= Days in year / Accounts receivable turnover
= 365/5.04 = 72.42 days
N Inventory turnover = Cost of goods sold / Average inventory *
= $740/$185 = 4
*Average inventory = (Inventory Year 1 + Inventory Year 2) / 2
= ($170+$200)/2 = $185
O Average sale period = Days in year / Inventory turnover
= 365/4 = 91.25 days
P Times interest eraned ratio = Earnings before interest and taxes / Interest Expenses
= $125/20 = 6.25 times
Q dEbt equity ratio = Total Liabilities / Stockholders' eQqity
= $1395/975
1.430769231